# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v1

@@ B.3 Proof of Theorem 5
$\left|v_{\gamma}^{\prime}\right|_{\mathrm{span}}=\left|v_{\gamma}^{*}\right|_{\mathrm{span}}\leq\frac{|S|}{\delta^{\prime}}$
which is uniform in $\gamma$. This and Theorem 3 implies Theorem 5. ∎
Appendix C Proofs for Section 5
C.1 Proof of Proposition 5.1
@@ Appendix C Proofs for Section 5
$0$ $\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (C.1)
@@
$\sum_{t=0}^{n-1}[\alpha_{\kappa}^{*}-r(X_{t},A_{t})]=\widetilde{O}(\left|h_{\kappa}^{*}\right|_{\mathrm{span}}\sqrt{n}).$
This implies that
@@
$0$ $\leq\alpha^{*}_{\kappa}-\overline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (C.2)
@@
Since $\pi_{\mathrm{RL}}$ only uses deterministic actions and $\left\{{\delta_{a}:a\in A}\right\}\subset\mathcal{Q}$, $\pi_{\mathrm{RL}}\in\Pi_{\mathrm{H}}(\mathcal{Q})$. Therefore, going back to (C.1), we have that
@@
Finally, to conclude the proposition, we note that if $\underline{\alpha}$ is replaced by $\overline{\alpha}$, the derivation in (C.1) is still valid. This, coupled with (C.2), yields the limsup version of Theorem 5.1. ∎
C.2 Proof of Theorem 6
@@
Also, by Markov optimality in classical MDPs [6], $\sup_{\pi\in\Pi_{\mathrm{S}}}\underline{\alpha}(\mu,\pi,\kappa)=\sup_{\pi\in\Pi_{\mathrm{H}}}\underline{\alpha}(\mu,\pi,\kappa)$. Since $\epsilon$ can be arbitrarily small, these inequalities imply the liminf version of Theorem 6. The same argument holds when $\underline{\alpha}$ is replaced with $\overline{\alpha}$.
To show the second claim, we note that the same argument as in the proof of Theorem 3 implies that if $\left|v_{\gamma}^{\prime}\right|_{\mathrm{span}}$ is uniformly bounded in $\gamma$, then (3.2) has a solution $(u^{\prime},\alpha^{\prime})$. Moreover, the argument for Theorem 1 will imply that $\alpha^{\prime}$ is the optimal average-reward
@@
So, we only need to show that $\left|v_{\gamma}^{\prime}\right|_{\mathrm{span}}$ is uniformly bounded in $\gamma$. This is achieved by employing the same proof as Theorem 4. As the proof idea is already used in both Theorem 4 and 5, we omit further detail. ∎
C.3 Proof of Corollary 6.1
@@
By Theorem 4 and 6, solutions $(u^{*},\alpha^{*})$ and $(u^{\prime},\alpha^{\prime})$ to (3.1) and (3.2) exists under the assumptions of Corollary 6.1. Hence, by Theorem 1 and 6,

### v2

@@ Appendix D Proof of Theorem 6
Then, we split $\left|v^{\prime}_{\gamma}\right|_{\mathrm{span}}$ as follows:
$\left|v_{\gamma}^{\prime}\right|_{\mathrm{span}}$ $=v_{\gamma}^{\prime}(s_{\vee})-v_{\gamma}^{\prime}(s_{\wedge})$ (D.3)
$\leq\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(s_{\vee})-\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(s_{\wedge})+\epsilon$
$\leq v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(s_{\vee})-\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(s_{\wedge})+2\epsilon$
$\leq\underbrace{v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(s_{\vee})-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})}_{\xi_{2}}+\underbrace{v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})-v_{\gamma}^{\pi,p_{\epsilon}}(s_{\wedge})}_{\xi_{1}}+2\epsilon$
for any $\pi\in\Pi_{\mathrm{H}}$.
Upper-Bounding $\xi_{1}$
Following the same construction as in the proof of Theorem 5 (but swapping the roles of controller and adversary), we bound $\xi_{1}$ by considering a two-phase history-dependent control policy as follows.
Fix $y\in C_{p_{k}}$ where $p_{\epsilon}\in G_{p_{k}}$. By Lemma 8 there exists a stationary controller policy $\eta$ such that
$\max_{w\in S}E^{\eta,p_{\epsilon}}_{w}\tau_{y}\leq\frac{|S|}{\delta^{\prime}}$
Thus, define a history-dependent two-phase controller $\pi=(\pi_{0},\pi_{1},\dots)$ by
$\pi_{t}(\cdot|h_{t})=\begin{cases}\eta(\cdot|s_{t}),&\text{if }s_{0},\dots,s_{t}\neq y,\\[2.0pt] \Delta_{\epsilon}(\cdot|s_{t}),&\text{otherwise.}\end{cases}$
Then, we have that for all $x\in S$
$v_{\gamma}^{\pi,p_{\epsilon}}(x)$ $=E^{\pi,p_{\epsilon}}_{x}\sum_{k=0}^{\tau_{y}-1}\gamma^{k}r(X_{k},A_{k})+E^{\pi,p_{\epsilon}}_{x}\gamma^{\tau_{y}}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ (D.4)
$\geq E^{\pi,p_{\epsilon}}_{x}\gamma^{\tau_{y}}E^{\pi,p_{\epsilon}}_{x}\left[\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})\middle|\mathcal{H}_{\tau_{y}}\right]$
$=v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)E^{\pi,p_{\epsilon}}_{x}\gamma^{\tau_{y}}.$
where the last equality follows from the same argument as in (5.9). Therefore, we have that
$v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)-v_{\gamma}^{\pi,p_{\epsilon}}(x)$ $\leq v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)E^{\pi,p_{\epsilon}}_{x}\gamma^{\tau_{y}}$
$=(1-\gamma)v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)E^{\pi,p_{\epsilon}}_{x}\frac{1-\gamma^{\tau_{y}}}{1-\gamma}$
$\leq E^{\pi,p_{\epsilon}}_{x}\tau_{y}$
where the last inequality follows from $0\leq v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)\leq 1/(1-\gamma)$ and $(1-\gamma^{t})/(1-\gamma)=\sum_{k=0}^{t-1}\gamma^{k}\leq t$.
On the other hand, mirroring (5.8), we have that $E^{\pi,p_{\epsilon}}_{x}\tau_{y}=E_{x}^{\eta,p_{\epsilon}}\tau_{y}$. Therefore, choosing $x=s_{\wedge}$, $y=y_{\vee}\in C_{p_{k}}$, and $\eta=\eta^{y_{\vee},p_{\epsilon}}$ given by Lemma 8, we conclude that
$\xi_{1}\leq E^{\pi,p_{\epsilon}}_{s_{\wedge}}\tau_{y_{\vee}}=E^{\eta,p_{\epsilon}}_{s_{\wedge}}\tau_{y_{\vee}}\leq\frac{|S|}{\delta^{\prime}}.$
Upper-Bounding $\xi_{2}$
Let $T_{k}:=\inf\left\{{t\geq 0:X_{t}\in C_{p_{k}}}\right\}$. To bound $\xi_{2}$, note that
$v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(s_{\vee})-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})$ $\leq E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=0}^{T_{k}-1}\gamma^{t}r(X_{t},A_{t})+E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=T_{k}}^{\infty}\gamma^{t}r(X_{t},A_{t})-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})$
$\stackrel{{\scriptstyle(i)}}{{=}}E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=0}^{T_{k}-1}\gamma^{t}r(X_{t},A_{t})+E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}E_{X_{T_{k}}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=0}^{\infty}\gamma^{t}r(X_{t},A_{t})-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})$
$\stackrel{{\scriptstyle(ii)}}{{=}}E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=0}^{T_{k}-1}\gamma^{t}r(X_{t},A_{t})+E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(X_{T_{k}})-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y_{\vee})$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}\sum_{t=0}^{T_{k}-1}\gamma^{t}r(X_{t},A_{t})$
$\leq E_{s_{\vee}}^{\Delta_{\epsilon},p_{\epsilon}}T_{k},$
where $(i)$ applies the strong Markov property, $(ii)$ recalls the definition of $v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}$, and $(iii)$ follows from $y_{\vee}\in C_{p_{k}}$ achieving the argmax.
Then, following the same argument as in Theorem 5 and the construction of $K_{p}$, we conclude that
$\xi_{2}\leq\max_{k\leq|B^{\prime}|}\sup_{p\in\overline{G_{p_{k}}},\;\Delta:S\rightarrow\mathcal{Q}}E_{s_{\vee}}^{\Delta,p}T_{k}\leq 1+\sum_{k\leq|B^{\prime}|}\sup_{\Delta:S\rightarrow\mathcal{Q}}e^{\top}(I-M_{\Delta,p_{k}}^{p_{k}})e<\infty.$
Remark 6.
If the adversary is communicating, then $C_{p}=S$. In this case, Lemma 8 holds with $y\in S$ arbitrary, and the term corresponding to transient states is vacuous, and $\xi_{2}\leq 0$. Thus, the compactness of $\mathcal{Q}$ is not required.
Combining the bounds for $\xi_{1}$ and $\xi_{2}$, we have shown that $\left|v^{\prime}_{\gamma}\right|_{\mathrm{span}}$ is uniformly bounded. By the same argument for Theorem 4, the uniform boundedness of $\left|v^{\prime}_{\gamma}\right|_{\mathrm{span}}$ implies the existence of $(u^{\prime},\alpha^{\prime})$ solving the average-reward inf-sup Bellman equation (3.2). ∎
Appendix E Proof of Corollary 7.3
We show the two claims of Corollary 7.3 by discussing how the overlap-connectedness assumptions are related to weak communication.
Lemma 9 (Overlap-Connected Controller).
If $\mathcal{R}_{\Delta}$ is overlap-connected and all $\mathcal{P}_{s}$, $s\in S$ are convex, then $\Delta$ is weakly communicating with $C_{\Delta}=\bigcup_{R\in\mathcal{R}_{\Delta}}R.$
Proof of Lemma 9.
Let $C_{\Delta}=\bigcup_{R\in\mathcal{R}_{\Delta}}R$. By the definition of $\mathcal{R}_{\Delta}$, for each $R\in\mathcal{R}_{\Delta}$, there exists $p^{(R)}$ such that $p^{(R)}_{\Delta}$ has $R$ as one of its closed recurrent classes.
Let $\mathcal{R}_{\Delta}(s):=\left\{{R\in\mathcal{R}_{\Delta}:s\in R}\right\}$. Then, define
[…]
By S-rectangularity and convexity, $q\in\mathcal{P}$.
By the overlap-connectedness of $\mathcal{R}_{\Delta}$, for all $s,s^{\prime}\in C_{\Delta}$ there exists $R_{0},\dots,R_{k}$ such that $s\in R_{0}$, $s^{\prime}\in R_{k}$, and $R_{i}\cap R_{i+1}\neq\varnothing$ for all $0\leq i\leq k-1$. Then, pick an arbitrary element $y_{i+1}\in R_{i}\cap R_{i+1}$.
Let $y_{0}=s$ and $y_{k+1}=s^{\prime}$. As for all $i$, $y_{i}$ and $y_{i+1}$ are contained in $R_{i}$, there exists a path of positive probability under $p^{(R_{i})}_{\Delta}$ that connects $y_{i}$ and $y_{i+1}$. By the definition of $q$, $q_{\Delta}(\cdot|s)\geq\frac{1}{|\mathcal{R}_{\Delta}(s)|}p_{\Delta}^{(R_{i})}(\cdot|s)$ for all $s\in C_{\Delta}$. It follows that the path leads $y_{i}$ to $y_{i+1}$ has positive probability under $q$ for each $i$. Therefore, $s$ leads to $s^{\prime}$ under $q_{\Delta}$; i.e. there exists $N\geq 1$ s.t. $q_{\Delta}^{N}(s^{\prime}|s)>0$.
To verify weak communication, we also need to check that any $z\in C_{\Delta}^{c}$ is transient under all $p\in\mathcal{P}$. This is straightforward as, by definition, $z\in C_{\Delta}^{c}$ implies that $z\notin R$ for any $R\in\mathcal{R}_{\Delta}$; i.e. it is transient under all $p_{\Delta}$, $p\in\mathcal{P}$. This proves Lemma 9. ∎
Similarly, we also show that if $\mathcal{Q}$ is convex, an overlap-connected $\mathcal{R}_{p}$ will imply the weak communication of $p\in\mathcal{P}$.
Lemma 10 (Overlap-Connected Controller).
If $\mathcal{R}_{p}$ is overlap-connected and $\mathcal{Q}$ is convex, then $p$ is weakly communicating with $C_{p}=\bigcup_{R\in\mathcal{R}_{p}}R.$
Proof of Lemma 10.
. Let $C_{p}=\bigcup_{R\in\mathcal{R}_{p}}R$. By the definition of $\mathcal{R}_{p}$, for each $R\in\mathcal{R}_{p}$, there exists $\Delta_{R}$ such that $p_{\Delta_{R}}$ has $R$ as one of its closed recurrent classes.
Let $\mathcal{R}_{p}(s):=\left\{{R\in\mathcal{R}_{p}:s\in R}\right\}$. Then, define
$\eta(\cdot|s):=\begin{cases}\displaystyle\frac{1}{|\mathcal{R}_{p}(s)|}\sum_{R\in\mathcal{R}_{p}(s)}\Delta_{R}(\cdot|s),&\text{if }s\in C_{p},\\ \displaystyle\text{any }\mu\in\mathcal{Q},&\text{otherwise.}\end{cases}$
By convexity of $\mathcal{Q}$, $\eta\in\left\{{S\rightarrow\mathcal{Q}}\right\}$.
By the overlap-connectedness of $\mathcal{R}_{p}$, for all $s,s^{\prime}\in C_{p}$ there exists $R_{0},\dots,R_{k}$ such that $s\in R_{0}$, $s^{\prime}\in R_{k}$, and $R_{i}\cap R_{i+1}\neq\varnothing$ for all $0\leq i\leq k-1$. Then, pick an arbitrary element $y_{i+1}\in R_{i}\cap R_{i+1}$.
Let $y_{0}=s$ and $y_{k+1}=s^{\prime}$. As for all $i$, $y_{i}$ and $y_{i+1}$ are contained in $R_{i}$, there exists a path of positive probability under $p_{\Delta_{R_{i}}}$ that connects $y_{i}$ and $y_{i+1}$. By the definition of $\eta$,
$p_{\eta}(\cdot|s)=\sum_{a\in A}\eta(a|s)p(\cdot|s,a)\geq\frac{1}{|\mathcal{R}_{p}(s)|}\sum_{a\in A}\Delta_{R_{i}}(a|s)p(\cdot|s,a)=\frac{1}{|\mathcal{R}_{p}(s)|}p_{\Delta_{\mathbb{R}_{i}}}(\cdot|s)$
for all $s\in C_{p}$. It follows that the path leads $y_{i}$ to $y_{i+1}$ has positive probability under $p_{\eta}$ for each $i$. Therefore, $s$ leads to $s^{\prime}$ under $p_{\eta}$; i.e. there exists $N\geq 1$ s.t. $p_{\eta}^{N}(s^{\prime}|s)>0$.
To verify weak communication, we also need to check that any $z\in C_{p}^{c}$ is transient under all $p\in\mathcal{P}$. This is straightforward as, by definition, $z\in C_{p}^{c}$ implies that $z\notin R$ for any $R\in\mathcal{R}_{p}$; i.e. it is transient under all $p_{\Delta}$, $\Delta:S\rightarrow\mathcal{Q}$. This proves Lemma 10. ∎
Proof of Corollary 7.3.
By Lemma 9, the convexity and overlap-connectedness assumption in the first statement of Corollary 7.3 ensures that every $\Delta:S\to\mathcal{Q}$ is weakly communicating. Combined with the compactness of $\mathcal{Q}$ and of each $\mathcal{P}_{s}$ for $s\in S$, the assumptions of Theorem 5 are satisfied. Hence, (3.1) admits a solution.
Similarly, by Lemma 10, the convexity and overlap-connectedness assumption in the latter statement implies that every $p\in\mathcal{P}$ is weakly communicating. Together with the compactness of $\mathcal{Q}$ and of each $\mathcal{P}_{s}$, this verifies the assumptions of Theorem 6. Therefore, (3.2) admits a solution. ∎
Appendix F Proofs for Section 6
F.1 Proof of Proposition 6.1
@@ Appendix F Proofs for Section 6
$0$ $\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (F.1)
@@
$\sum_{t=0}^{n-1}[\alpha_{\kappa}^{*}-r(X_{t},A_{t})]=\widetilde{O}(\left|h_{\kappa}^{*}\right|_{\mathrm{span}}\sqrt{n})$
[…]
@@
$0$ $\leq\alpha^{*}_{\kappa}-\overline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (F.2)
@@
Since $\pi_{\mathrm{RL}}$ only uses deterministic actions and $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, $\pi_{\mathrm{RL}}\in\Pi_{\mathrm{H}}(\mathcal{Q})$. Therefore, going back to (F.1), we have that
@@
Finally, to conclude the proposition, we note that if $\underline{\alpha}$ is replaced by $\overline{\alpha}$, the derivation in (F.1) is still valid. This, coupled with (F.2), yields the limsup version of Theorem 6.1. ∎
F.2 Proof of Theorem 8
@@
Also, by Markov optimality in classical MDPs [18], $\sup_{\pi\in\Pi_{\mathrm{S}}}\underline{\alpha}(\mu,\pi,\kappa)=\sup_{\pi\in\Pi_{\mathrm{H}}}\underline{\alpha}(\mu,\pi,\kappa)$. Since $\epsilon$ can be arbitrarily small, these inequalities imply the liminf version of the first claim Theorem 8. The same argument holds when $\underline{\alpha}$ is replaced with $\overline{\alpha}$.
To show the second claim, we note that the same argument as in the proof of Theorem 1 will imply that the solution $\alpha^{\prime}$ to (3.2) is the optimal average-reward
@@
So, $\alpha^{\prime}$ corresponds to the inf-sup control value, which is shown to be equal to $\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})$. The same holds true if $\underline{\alpha}$ is replaced by $\overline{\alpha}$. ∎
F.3 Proof of Corollary 8.1
@@
By Theorem 5 and 6, solutions $(u^{*},\alpha^{*})$ and $(u^{\prime},\alpha^{\prime})$ to (3.1) and (3.2) exists under the assumptions of Corollary 8.1. Hence, by Theorem 1 and 8,